f(y)= y + 3/y^2-3y
what is the domain function
answer i interval notation

Answer by lwsshak3(11628)

f(y)= y + 3/y^2-3y
what is the domain function
answer i interval notation
**
you can see that if the denominator=0, f(y) becomes undefined
so, set denominator=0, then solve for values of y that make the denominator=0
y^2-3y=0
y(y-3)=0
y≠0 and y≠3
domain: (-∞,0) U (0,3) U (3,∞)